The owner of a small convenience store is trying to decide whether to discontinue selling magazines. He suspects that only 6% of the customers buy a magazine and thinks that he might be able to use the display space to sell something more profitable. Before making a final decision, he decides that for one day he’ll keep track of the number of customers and whether or not they buy a magazine. Assuming the owner is correct in thinking that 6% of the customers purchase magazines, how many customers should he expect before someone buys a magazine? And is this problem Geometric or Binomial?

Answer choices

6 customers - Binomial

16 customers - Geometric

6 customers - Geometric

16 customers - Binomial
Answer #1

Number of customers expected = 1/p = 1/0.06 = 16

Hence,

16 customers - Geometric

Option B is correct.